OpenStudy (anonymous):

Completely factor the expression: 112x^2-567y^2 Could you show me the prime factorizations? What do they both go into, if I can figure this out I can do the rest, so please do not tell me the answer yet.

OpenStudy (anonymous):

What I have so far is they both go into 7 112=7*16 567= 7*81

OpenStudy (anonymous):

If this is the right combination, then how do I set that up, I know I know how to do this it is just not coming to me. would it be (7x-16y)(7x-81y)

OpenStudy (anonymous):

\[7(16x^2 - 81y^2) = 7(4x + 9y)(4x-9y)\]

OpenStudy (anonymous):

That is the answer? Can you show your work.

OpenStudy (anonymous):

Why can it not be wrote like this7(16x-81y)^2 ?

OpenStudy (anonymous):

Because of the binomial theorem. \[(a-b)*(a+b) = a^2 -a*b + a*b - b^2\]

OpenStudy (anonymous):

How do you work it out with this theorem? what is a and what is b?

OpenStudy (anonymous):

a and b are just variables. You did a good job factoring out the common factor of 7. After that, you factor what is called a difference of squares. 16 is 4 squared and 81 is 9 squared. 4 = a and 9 = b in this case. If you plug in those numbers you will see that my formula matches your question.

OpenStudy (anonymous):

Ok, lets do it, give me a minute.

OpenStudy (anonymous):

(a-b)*(a+b)=a^2-a * b+a*b-b^2 (4-9) * (4+9)=4^2-4 *9+4*9-9^2 -5*13=16-4*9+4*9- 81 -65=12*9 -65=108+4*9-81 -65=112*9-81 -65=1008-81 -65=927 This is what I have, but it doesnt seem right. I used PEMDAS to calculate it.

OpenStudy (anonymous):

wat did you do.... 16 - 81 = -5*13.... -4*9 + 4*9 = 0

OpenStudy (anonymous):

PEMDAS is ORDER of operations. MEANING you do MULTIPLICATION and DIVISION BEFORE BEFORE BEFORE addition and subtraction. What grade are you in?

OpenStudy (anonymous):

I did it in order, It wasn't coming out = so I didn't think it was right. I did the exponents first. Well you dont have to be so mean about it!

OpenStudy (anonymous):

(a-b)*(a+b)=a^2-a * b+a*b-b^2 (4-9) * (4+9)=4^2-4 *9+4*9-9^2 -5*13=16-4*9+4*9-81 -65=16-|-144+36-81 -65=160+36-81 -65=196-81 -65=115 It still did not= out. If you are going to have an attitude towards helping me, you don't have to I will repost it and get help from someone else!

OpenStudy (anonymous):

\[(4-9)*(4+9) =(4^2) + (- 4*9) + (4*9) - (9^2) = -65 = -5*13\] I'm using Capitals to emphasize my point. And I was seriously asking what grade you are in, I'm a Sophomore in College, I want to know how much math you have taken.